A race car driver is driving his 2000kg racecar through a turn at 60m/s. The radius of the curve is 200m.

a. What is the minimum force of friction so that he does not slide out into the wall?

b. If this is a flat track what is the minimum coefficient of friction between the tires and the road?
1
Expert's answer
2020-11-20T07:15:01-0500

a)


"F_f=\frac{mv^2}{r}=\frac{(2000)(60)^2}{200}=36000\ N"

b)


"\mu=\frac{F_f}{mg}=\frac{36000}{2000\cdot9.8}=1.84"